# Niobium coaxial cavities with internal quality factors exceeding 1.5 billion for circuit quantum electrodynamics

### v1

@@ Appendix E Nb oxide thickness analysis
Where the ratio of the volume densities of Nb atoms in metal to different oxide phases () are listed in Fig. S5. The inelastic mean free path (IMFP) values, , for Nb2O5, NbO2, and NbO can be found in Ref. Powell and Jabolonski (2011). These values are specified normal to the surface (= 90) of the Kratos Axis Nova XPS instrument, which has an angle of 45 between the X-ray source and the axis of the electron lens. For direct determination, a fine sputter etch (15 s per cycle) was used, with the cycle repeated until the relative concentrations of both niobium and oxygen reached a constant value. Fig. S6 (a) shows the niobium and oxygen concentrations from the survey scans as a function of sputter cycle. By 11 cycles the 3 week air exposed cavity show stable concentrations of oxygen and niobium, while the cavity exposed to air for 30 minutes reaches a stable value in 8 sputter cycles. Using the calibration in Appendix. D, this gives a directly measured oxide thickness of and for 3 week and 30 minute air exposure respectively. The Nb oxide thickness extracted via the indirect method was determined to be and for 3 weeks and 30 minutes respectively. The estimated oxide thickness from the Nb thin film used for sputter calibration in Appendix. D, and Nb cavities with different air exposure duration are shown in the Fig. S6 (b) table, along with the relative thicknesses of the three Nb oxidation states (NbO, NbO2, and Nb2O5).

### v2

@@ Appendix E Nb oxide thickness analysis
Where the ratio of the volume densities of Nb atoms in metal to different oxide phases () are listed in Table. 2. The inelastic mean free path (IMFP) values, , for Nb2O5, NbO2, and NbO can be found in Ref. Powell and Jabolonski (2011). These values are specified normal to the surface (= 90∘) of the Kratos Axis Nova XPS instrument, which has an angle of 45∘ between the X-ray source and the axis of the electron lens. For direct determination, a fine sputter etch (15 s per cycle) was used, with the cycle repeated until the relative concentrations of both niobium and oxygen reached a constant value. Fig. S5 shows the niobium and oxygen concentrations from the survey scans as a function of sputter cycle. By 11 cycles the 3 week air exposed cavity show stable concentrations of oxygen and niobium, while the cavity exposed to air for 30 minutes reaches a stable value in 8 sputter cycles. Using the calibration in Appendix. D, this gives a directly measured oxide thickness of and for 3 week and 30 minute air exposure respectively. The Nb oxide thickness extracted via the indirect method was determined to be and for 3 weeks and 30 minutes respectively. The estimated oxide thickness from the Nb thin film used for sputter calibration in Appendix. D, and Nb cavities with different air exposure duration are shown in Table E, along with the relative thicknesses of the three Nb oxidation states (NbO, NbO2, and Nb2O5).
| Species | Nb 75nm Thin Film | Nb 2-4 | |
| Nb 2-4 | (+3 wks) | ||
| (+30 min) | |||
| Nb2O5 | 4.50.01 nm | 4.40.01 nm | 2.40.04 nm |
| NbO2 | 2.10.02 nm | 0.40.01 nm | 0.50.02 nm |
| NbO | 0.50.02 nm | 0.10.03 nm | 0.60.01 nm |
| Total | 7.10.02 nm | 4.90.01 nm | 3.50.01 nm |